# Superposed Decoding: Multiple Generations from a Single Autoregressive Inference Pass

### v2

@@ Abstract
Many applications today provide users with multiple auto-complete drafts as they type, including GitHub’s code completion, Gmail’s smart compose, and Apple’s messaging auto-suggestions. Under the hood, language models support this by running an autoregressive inference pass to provide a draft. Consequently, providing $k$ drafts to the user requires running an expensive language model $k$ times. To alleviate the computation cost of running $k$ inference passes, we propose Superposed Decoding, a new decoding algorithm that generates $k$ drafts at the computation cost of one autoregressive inference pass. We achieve this by feeding a superposition of the most recent token embeddings from the $k$ drafts as input to the next decoding step of the language model. At every inference step we combine the $k$ drafts with the top-$k$ tokens to get $k^{2}$ new drafts and cache the $k$ most likely options, using an n-gram interpolation with minimal compute overhead to filter out incoherent generations. Our experiments show that $k$ drafts from Superposed Decoding are at least as coherent and factual as Nucleus Sampling and Greedy Decoding respectively, while being at least 2.44×2.$4$ faster for $k\geq 3$. In a compute-normalized setting, user evaluations demonstrably favor text generated by Superposed Decoding over Nucleus Sampling. Code and more examples open-sourced at https://github.com/RAIVNLab/SuperposedDecoding.
@@ 1 Introduction
Our experiments demonstrate that Superposed Decoding on Llama-2-7B [40] can generate $k\geq 1$ drafts that are as coherent, in terms of perplexity, as Nucleus Sampling (Section 4.1). However, SPD achieves this with a single inference pass, making it at least 2.44×2.$4$ faster than any other standard decoding methods for $k\geq 3$ (Section 4.3). The ability to generate multiple drafts at the cost of one also increases the coverage significantly for fact-based evaluation tasks like TriviaQA and Natural questions – where SPD increases the chance of generating the correct answer by at least $5\%$ when using 3 drafts (Section 4.2). Finally, through an extensive human evaluation for a wide range of prefixes, we show that SPD generations are as preferred as Nucleus Sampling in direct comparison (1v1) while outperforming by up to $20\%$ in a compute normalized setting (3v1 and 3v2) (Section 4.4).
@@ 2 Related Work
Superposed Decoding (SPD), on the other hand, can be easily integrated out of the box without any additional training on any language model. Further, Superposed Decoding is complementary to Medusa and multi-token prediction, as well as other efficiency techniques like speculative decoding [29, 25, 5], quantization [9, 31, 18], pruning [12, 38, 27], and architectural optimizations [36, 2, 1, 42, 26, 10].
@@ 3.3 Superposed Decoding Semantically Approximates Beam Search
Figure 4: Qualitative text generations in a compute-normalized setting for Superposed Decoding and Nucleus Sampling with prefixes sampled from OpenWebText. See Appendix G for more.
@@ 4 Experiments
We evaluate Superposed Decoding by analyzing generation quality, factuality, and latency. We demonstrate that Superposed Decoding improves over Nucleus Sampling and Greedy Decoding in generation perplexity (Section 4.1), fact-based benchmarks (Section 4.2), and wall-clock time (Section 4.3). Finally, we conduct a user study highlighting that users prefer Superposed Decoding over Nucleus Sampling in a compute-normalized setting (Section 4.4). We also include qualitative examples of Superposed Decoding generations in Figure 4, with more in Appendix G.
We implement Superposed Decoding on the base pre-trained version of Llama-2-7B [41] for the majority of our experiments, running on a single A40 GPU. We do not retrain or fine-tune any part of the model. For perplexity evaluations, we use Llama-2-70B over eight A40 GPUs.
@@ 4.3 Latency
Superposed Decoding presents a significant theoretical reduction in latency, but it is important to investigate how well these savings translate to real-world settings. In Figure 7, we show that the only additional cost incurred by Superposed Decoding is due to dictionary lookups, barring which Superposed Decoding has near-constant compute cost. Even so, the total cost of Superposed Decoding is significantly lower than other decoding methods, with Superposed Decoding $\mathbf{2.44}\times$ faster on three drafts and 3.54$\times$ faster on eight compared to Nucleus Sampling, the second fastest.
@@ 6 Discussion and Conclusion
While we demonstrate that Superposed Decoding is an effective technique for multiple draft generation, Superposed Decoding is limited by the quality of the n-gram models used, which are essential for maintaining coherence. In addition, while Superposed Decoding drafts are syntactically diverse, they are not often semantically diverse. We suspect that the orthogonality of token embeddings discovered by Jiang et al. [22] is a potential solution. While our initial experiments did not show diversity gains, we believe that orthogonality is promising and is a logical next step for future work. We also note that mechanisms, like batching, to increase throughput of decoding algorithms are complementary to Superposed Decoding.
In conclusion, we present Superposed Decoding, a novel decoding method that superposes token embeddings to generate multiple short generations at the cost of one. We demonstrate that Superposed Decoding improves on Nucleus Sampling in terms of generation quality and human preference. The plurality of choices from Superposed Decoding also leads to better performance on common benchmarks. We envision that the latency reduction from Superposed Decoding will make it practical to apply large pre-trained language models on drafting tasks where compute is often a barrier for deployability. Finally, Superposed Decoding can be deployed in messaging applications using n-grams personalized to each user, where the number of n-grams can be reduced to save storage without compromising performance.
@@ Appendix D Standard Deviations for Perplexity Evaluation
We calculate the standard deviations across the 5000 OpenWebText prefixes used for evaluation for Llama-2-7B and Mistral 7B and show them below in Tables 4 and 5. We note that Superposed Decoding has higher standard deviation than Nucleus Sampling and Beam Search when implemented on Llama-2-7B but that the standard deviations equalize on Mistral 7B. Despite this high variance, human evaluation results in Section 4.4 affirm that Superposed Decoding is consistently competitive in coherency.

### v4

@@ Abstract
Many applications today provide users with multiple auto-complete drafts as they type, including GitHub’s code completion, Gmail’s smart compose, and Apple’s messaging auto-suggestions. Under the hood, language models support this by running an autoregressive inference pass to provide a draft. Consequently, providing $k$ drafts to the user requires running an expensive language model $k$ times. To alleviate the computation cost of running $k$ inference passes, we propose Superposed Decoding, a new decoding algorithm that generates $k$ drafts at the computation cost of one autoregressive inference pass. We achieve this by feeding a superposition of the most recent token embeddings from the $k$ drafts as input to the next decoding step of the language model. At every inference step we combine the $k$ drafts with the top-$k$ tokens to get $k^{2}$ new drafts and cache the $k$ most likely options, using an n-gram interpolation with minimal compute overhead to filter out incoherent generations. Our experiments show that $k$ drafts from Superposed Decoding are at least as coherent and factual as Nucleus Sampling and Greedy Decoding respectively, while being at least 2.44×2.$4$ faster for $k\geq 3$. In a compute-normalized setting, user evaluations demonstrably favor text generated by Superposed Decoding over Nucleus Sampling. Superposed Decoding can also be combined with other decoding strategies, resulting in universal coverage gains when scaling inference time compute. Code and more examples open-sourced at https://github.com/RAIVNLab/SuperposedDecoding.
@@ 1 Introduction
Our experiments demonstrate that Superposed Decoding on Llama-2-7B [41] can generate $k\geq 1$ drafts that are as coherent, in terms of perplexity, as Nucleus Sampling (Section 4.1). However, SPD achieves this with a single inference pass, making it at least 2.44×2.$4$ faster than any other standard decoding methods for $k\geq 3$ (Section 4.3). The ability to generate multiple drafts at the cost of one also increases the coverage significantly for fact-based evaluation tasks like TriviaQA and Natural questions – where SPD increases the chance of generating the correct answer by at least $5\%$ when using 3 drafts (Section 4.2). Through an extensive human evaluation for a wide range of prefixes, we show that SPD generations are as preferred as Nucleus Sampling in direct comparison (1v1) while outperforming by up to $20\%$ in a compute normalized setting (3v1 and 3v2) (Section 4.4). Finally, we find that combining Superposed Decoding and other decoding strategies (e.g. Nucleus Sampling) results in substantial accuracy improvements when scaling inference compute (Section 4.5).
@@ 2 Related Work
Superposed Decoding (SPD), on the other hand, can be easily integrated out of the box without any additional training on any language model. Further, Superposed Decoding is complementary to other decoding methods like Medusa and multi-token prediction, as well as efficiency techniques like speculative decoding [30, 26, 6], quantization [10, 32, 19], pruning [13, 39, 28], and architectural optimizations [37, 2, 1, 43, 27, 11].
@@ 3.3 Superposed Decoding Semantically Approximates Beam Search
Figure 4: Qualitative text generations in a compute-normalized setting for Superposed Decoding and Nucleus Sampling with prefixes sampled from OpenWebText. See Appendix G.1 for more.
@@ 4 Experiments
We evaluate Superposed Decoding by analyzing generation quality, factuality, and latency. We demonstrate that Superposed Decoding improves over Nucleus Sampling and Greedy Decoding in generation perplexity (Section 4.1), fact-based benchmarks (Section 4.2), and wall-clock time (Section 4.3). We also conduct a user study highlighting that users prefer Superposed Decoding over Nucleus Sampling in a compute-normalized setting (Section 4.4). Finally, we show that Superposed Decoding improves performance while scaling inference compute (Section 4.5). We include example generations of Superposed Decoding in Figure 4, with more in Appendix G.1.
We implement Superposed Decoding on the base version of Llama-2-7B [42] for the majority of our experiments, running on one A40 GPU. We do not change model weights. For perplexity evaluations, we use Llama-2-70B on eight A40 GPUs. We assume a batch size of one for all experiments.
@@ 4.3 Latency
Superposed Decoding presents a significant theoretical reduction in latency, but it is important to investigate how well this translates to the real-world. In Figure 7, we show that dictionary lookups are the only additional cost incurred by Superposed Decoding, barring which Superposed Decoding has near-constant compute cost. Even so, the total cost of Superposed Decoding is significantly lower than other decoding methods, with Superposed Decoding $\mathbf{2.44}\times$ faster on three drafts and 3.54$\times$ faster on eight compared to Nucleus Sampling, the next fastest.
@@ 4.4 Human Evaluation
Table 2: Coverage on TriviaQA and Natural Questions in a normalized compute setting comparing vanilla Nucleus Sampling ($NS$) to the combination of Nucleus Sampling and Superposed Decoding ($NS_{SPDk}$), where $k$ is the number of Superposed Decoding drafts generated on top of each Nucleus Sampled generation. $NS_{SPD}$ results in better coverage at nearly every compute budget $n$.
Task Decoding Method Compute ($n$) 1 10 20 30 40 50 60 70 80 90 100 TriviaQA $NS_{~{}~{}~{}~{}~{}~{}~{}~{}}$ 51.04 68.75 70.31 71.87 72.92 74.48 74.74 75.26 75.78 76.30 76.56 $NS_{SPD2}$ 51.30 68.75 70.57 72.66 74.74 75.78 76.30 76.82 78.39 79.17 79.43 $NS_{SPD3}$ 51.82 70.57 74.22 75.52 77.34 77.87 78.39 78.65 79.17 79.43 79.95 Natural Questions $NS_{~{}~{}~{}~{}~{}~{}~{}~{}}$ 14.32 32.55 36.98 38.54 40.36 41.15 41.67 41.93 42.19 42.71 42.97 $NS_{SPD2}$ 15.36 31.25 34.90 38.02 39.84 41.41 41.67 42.45 43.75 43.75 43.75 $NS_{SPD3}$ 15.63 31.25 36.98 39.06 41.15 42.71 43.75 43.75 44.27 44.79 45.57
4.5 Inference-Time Scaling
Superposed Decoding also provides significant benefits for inference time compute scaling. Superposed Decoding is completely complimentary to other decoding methods, expanding semantic coverage at no extra compute. For instance, if Nucleus Sampling is used to generate $n$ drafts, Superposed Decoding with $k$ drafts can be spliced in at any timestep to strategically produce $nk$ drafts at no extra cost, using the Nucleus Sampled generations as prefixes. This bolsters every Nucleus Sampling generation with $k$ local searches (example in Appendix G.2).
This property is valuable for repeated sampling [3], a technique where the number of generations is scaled at inference time to increase coverage – the proportion of questions that can be answered correctly using at least one of the generations. While repeated sampling typically uses Nucleus Sampling, combining Superposed Decoding and Nucleus Sampling ($NS_{SPD}$) produces even larger coverage gains. In Table 2, we compare the coverage of vanilla Nucleus Sampling and $NS_{SPD}$ in an equal compute setting up to 100 Nucleus Sampling drafts. We find that $NS_{SPD}$ results in higher coverage across the board on both TriviaQA [24] and Natural Questions [29], highlighting Superposed Decoding as a powerful method to increase the impact of scaling inference compute.
@@ 6 Discussion and Conclusion
While we demonstrate that Superposed Decoding is an effective technique for multiple draft generation, Superposed Decoding is limited by the quality of the n-gram models used, which are essential for maintaining coherence. In addition, while Superposed Decoding drafts are syntactically diverse, they are not often semantically diverse. We suspect that the orthogonality of token embeddings discovered by Jiang et al. [23] is a potential solution. While our initial experiments did not show diversity gains, we believe that orthogonality is promising and is a logical next step for future work. We also note that mechanisms, like batching, that increase throughput of decoding algorithms are complementary to Superposed Decoding.
In conclusion, we present Superposed Decoding, a novel decoding method that superposes token embeddings to generate multiple short generations at the cost of one. We demonstrate that Superposed Decoding improves on Nucleus Sampling in terms of generation quality and human preference. The plurality of choices from Superposed Decoding also leads to better performance on common benchmarks and expands coverage at scale. We envision that the latency reduction from Superposed Decoding will make it practical to apply large pre-trained language models on drafting tasks where compute is often a barrier for deployability. Finally, Superposed Decoding can be deployed in messaging applications using n-grams personalized to each user, where the number of n-grams can be reduced to save storage without compromising performance.
@@ Appendix D Standard Deviations for Perplexity Evaluation
We calculate the standard deviations across the 5000 OpenWebText prefixes used for evaluation for Llama-2-7B and Mistral 7B and show them below in Tables 5 and 6. We note that Superposed Decoding has higher standard deviation than Nucleus Sampling and Beam Search when implemented on Llama-2-7B but that the standard deviations equalize on Mistral 7B. Despite this high variance, human evaluation results in Section 4.4 affirm that Superposed Decoding is consistently competitive in coherency.
@@ G.1 Superposed Decoding
G.2 Nucleus Sampling and Superposed Decoding
In Figure 15, we show an example generation that outputs three drafts using Nucleus Sampling for several timesteps, followed by Superposed Decoding with $k=3$. This results in nine total drafts at the compute of only three Nucleus Sampling drafts. The final Superposed Decoding drafts maintain the wide coverage offered by Nucleus Sampling and complement it with additional local coverage.
Figure 15: Example output generated by combining Nucleus Sampling and Superposed Decoding. Superposed Decoding generates three drafts per Nucleus Sampling sample, resulting in nine total drafts.